OpenStudy (anonymous):

Rene earns $8.50 per hour on days other than Sunday and twice that rate on Sundays. Last week she worked a total of 40 hours, including 8 hours on Sunday. What were her earnings for the week? 1. $340 2. $398 3. $408 4. $476

OpenStudy (anonymous):

c i found it

ganeshie8 (ganeshie8):

anum lets put it on paper. total earnings = 32 x 8.5 + 8 x 2 x 8.5 = 272 + 136 = $408 is it clear ?

OpenStudy (anonymous):

yes